A small probe is made of stainless steel and has a length of 164.1 mm when being used to monitor the temperature in a furnace.
When the probe is removed for maintenance and allowed to cool to 20.0°C the length of the probe is measured as 162.2 mm.
What is the temperature of the furnace? (to 3 s.f and in °C)
[αsteel = 16.0×10−6 K−1]
1
Expert's answer
2020-05-21T13:46:09-0400

Solution.

"L=164.1mm;"

"L_0=162.2mm;"

"T_0=20^oC=293K;"

"\alpha_L=16.0\sdot10^{-6}K^{-1};"

"\Delta T-?;"

The relative change in linear dimension, which can be considered as a relative deformation, can be recorded:"\epsilon=\dfrac{\Delta L}{L_0}=\alpha_L\sdot\Delta T\implies\Delta T=\dfrac{\Delta L}{L_0\alpha_L};"

"\Delta L=164.1mm-162.2mm=1.9mm;"

"\Delta T=\dfrac{1.9mm}{162.2mm\sdot 16.0\sdot 10^{-6} K^{-1}}=732K;"

"\Delta T=T-T_0; \implies T=\Delta T+T_0;"

"T=732K+293K=1025K=752^oC;"

Answer:"T=752^oC."